I need help on this problem.
Solve the equation.
5x^2 - 30x + 40 = 0

Answer by hamda(7)


5x^2 - 30x + 40 = 0
5(x^2 - 6x + 8) = 0
x^2 - 6x + 8 = 0
x^2 - 4x - 2x + 8 = 0
x(x - 4) - 2(x - 4) = 0
(x - 4)(x - 2) = 0
x = 2 , x = 4 (answer)